# NyayaRAG: Realistic Legal Judgment Prediction with RAG under the Indian Common Law System

### v2

@@ 7.1 Judgment Prediction Performance
In the multi-label setting, the best accuracy was observed for the CaseText + Statutes + Precedents pipeline. This comprehensive context provides the model with structured legal knowledge, improving generalization across different outcome labels. Conversely, the Facts Only pipeline performed worst overall, reaffirming that factual narratives alone, without legal context, are insufficient for reliably predicting legal outcomes. The poor performance of the Facts + Statutes + Precedents pipeline in the single-label setting suggests that factual sections might lack the interpretive cues that full case texts offer when combined with legal references.
@@ 7.2 Explanation Generation Quality
Table 5: Inter-Annotator Agreement (IAA) statistics for expert evaluation of generated legal explanations across different pipeline settings for both Single and Multi partitions.
@@ 8 Ablation Study: Understanding the Role of Legal Context Components
9 Conclusion and Future Scope
@@
Future directions include extending to hierarchical verdict structures, integrating symbolic or graph-based retrieval, modeling temporal precedent evolution, and leveraging human-in-the-loop mechanisms. NyayaRAG marks a step toward court-aligned, explainable legal AI and sets the foundation for future research in retrieval-enhanced legal systems within underrepresented jurisdictions.

### v3

@@ 7.1 Judgment Prediction Performance
In the multi-label setting, the best accuracy was observed for the CaseText + Statutes + Precedents pipeline. This comprehensive context provides the model with structured legal knowledge, improving generalization across different outcome labels. Conversely, the Facts Only pipeline performed worst overall, reaffirming that factual narratives alone, without legal context, are insufficient for reliably predicting legal outcomes. The poor performance of the Facts + Statutes + Precedents pipeline in the single-label setting suggests that factual sections lack the interpretive cues that full case texts offer when combined with legal references.
@@ 7.2 Explanation Generation Quality
Table 5: Inter-Annotator Agreement (IAA) statistics for expert evaluation of generated legal explanations across different pipeline settings for both Single and Multi partitions
Expert Evaluation:
@@ 8 Ablation Study: Understanding the Role of Legal Context Components
9 Conclusions and Future Scope
@@
Future directions include extending to hierarchical verdict structures, integrating symbolic or graph-based retrieval, modeling temporal precedent evolution, and leveraging human-in-the-loop mechanisms. NyayaRAG marks a step toward court-aligned, explainable legal AI and sets the foundation for future research in retrieval-enhanced legal systems within under-represented jurisdictions.